OpenStudy (anonymous):

I know this is a long shot, but I believe I have a genius out there. use long division to find the quotient of the following problem. (12x3 + 9x - 11x2 + 19) / (4x + 3) (the 3 and 2 in this problem are hypen 3 and 2)

OpenStudy (anonymous):

start with \[\frac{12x^3 + 9x - 11x^2 + 19}{4x + 3}\] \[=\frac{12x^3-11x^2+9x+19}{4x+3}\] then if you can factor the numerator you can cancel, if not, you have to divide using long division

OpenStudy (anonymous):

actually forget the part about factoring, since 3 does not divide 19 you are going to do long division for this one, which i find impossible to write here

OpenStudy (anonymous):

no it is not \(\frac{2(5x+7)}{4x+3}\)

OpenStudy (anonymous):

can you send it to me?

OpenStudy (anonymous):

the answer i can send you instantly the method you need pencil and paper for

OpenStudy (anonymous):

Would love both w/the hopes of learning how to do it, but right now any thing you can send would help.

OpenStudy (anonymous):

chris-j@comcast.net

OpenStudy (anonymous):

the answer is \[3x^2-5x+6+\frac{1}{4x+3}\]

OpenStudy (anonymous):

the method is to divide as you would with numbers

OpenStudy (anonymous):

use the chris email. hmmm, would love to know the long version.

OpenStudy (anonymous):



OpenStudy (anonymous):

ok, any more tips, I really need to know how to do this.

OpenStudy (anonymous):

that is the first step, it is just like dividing numbers

OpenStudy (anonymous):

think like this: \(4x\) goes in to \(12x^3\) \(3x^2\) times just as you would with whole numbers then put the \(3x^2\) up top, multiply \(3x^2\times (4x+3)=12x^3+9x^2\) put that underneath, then subtract